OpenStudy (anonymous):

For the following ellipses, indicate whether the major axis is the x or y axis... 36. x^2/16 + y^2/9 =1 37. x^2/4 + y^2/49 = 1 38. x^2+y^2/2=1 39. 4x^2+9y^2=36 40. 8x^2+2y^2=64

OpenStudy (anonymous):

holy, where do we start?

OpenStudy (anonymous):

i have no clue, This stuff is so HARD!

OpenStudy (anonymous):

i meant what problem... :) i can help you...

OpenStudy (anonymous):

All of them, lol

OpenStudy (mertsj):

If the big number is under the x, the major axis is horizontal. If the big number is under the y, the major axis is vertical.

OpenStudy (anonymous):

so number 36 is x?

OpenStudy (anonymous):

exactly what @Mertsj , said but remember you must get 1 side equal to 1 also...

OpenStudy (anonymous):

called standard form...

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

take #6... which has the bigger denominator?

OpenStudy (anonymous):

x^2 or y^2?

OpenStudy (anonymous):

what number?

OpenStudy (anonymous):

the number on the bottom of the fraction.

OpenStudy (anonymous):

i know that, i meant what number are you on? 36?

OpenStudy (mertsj):

36 is x 37 is y 38 is y 39 is x 40 is y

OpenStudy (anonymous):

\[\frac{x^2}{16}+\frac{y^2}{9}=1\] maybe that will help.. which has the bigger denominator?

OpenStudy (anonymous):

i guess you got all the answers already... cya!

OpenStudy (anonymous):

Thank both of yahl.